OpenStudy (anonymous):

What Is A Multipule Of 27

OpenStudy (anonymous):

\[1\times 27\] \[2\times 27\] \[3\times 27\] \[...\]

myininaya (myininaya):

27m where m is integer

OpenStudy (anonymous):

hello!

myininaya (myininaya):

hey

OpenStudy (anonymous):

i wonder if the question is for multiples or factors of 27

myininaya (myininaya):

i listed all the multiples of 27 :)

OpenStudy (anonymous):

you did indeed